# What Do You Need for Diverse Trajectory Composition in Diffusion Planning?

### v2

@@ Abstract
In planning, stitching is an ability of algorithms to piece together sub-trajectories of data they are trained on to generate new and diverse behaviours. While stitching has historically been a strength of offline reinforcement learning, recent generative behavioural cloning (BC) methods have also shown proficiency at stitching. However, the main factors behind this are poorly understood, hindering the development of new algorithms that can reliably stitch. Focusing on diffusion planners trained via imitation, we find two properties are key contributors for composition: shift equivariance and local receptiveness. We use these two properties to explain architecture, data, and inference choices in existing generative BC methods based on diffusion planning, including replanning frequency, data augmentation, and data scaling. Experimental comparisons show that (1) while locality is more important than shift equivariance in creating a diffusion planner capable of composition, both are crucial (2) enabling these properties through relatively simple architecture choices can be competitive with more computationally expensive methods such as replanning or scaling data, and (3) simple inpainting-based guidance can guide architecturally compositional models to enable generalization in goal-conditioned settings.
1 Introduction
The recent resurgence in generative diffusion models [58, 30] has inspired a new wave of work investigating the potential for these models to act as planners for robotic agents [34, 13]. One surprising ability these diffusion planning models have is compositionality: the ability to take component elements seen in samples of their dataset and compose them in new, kinematically feasible ways [20, 44].
One form of compositionality especially important in diffusion planning is stitching, where some diffusion planning methods are capable of composing sub-trajectories into new trajectories distinct from the training data. Historically, stitching has been seen as one of the primary benefits of temporal difference (TD) learning [38], both in offline and online RL. Planning with diffusion models thus potentially offers some of the stitching benefits of TD learning without needing to handle RL-specific challenges.
Despite their recent success in vision, when and why diffusion models compose in the planning domain is not well understood. Contrary to toy experiments in prior works [1, 34], the convolutional neural network (CNN)-based backbones commonly used in diffusion planners often struggle with composition [9], instead regressing to memorization of training examples [2].
This gap between the promise of diffusion planners to enable stitching and reality that they frequently struggle with composition is crucial to address, as diffusion offers an appealing alternative to other stitching approaches. RL suffers from many challenges when stitching such as the deadly triad [33, 22, 64] and slow temporal credit assignment [12], while non-generative BC stitching approaches require nettlesome data augmentation [26]. The capability to sample a diverse set of new trajectories at test time is important property for planners broadly. Reasons for this include being able to select solutions from a candidate set [69], encouraging multi-modal behaviour [40], or generating diverse data to train imitation learning or offline RL algorithms [38].
To address this gap, we analyze a variety of approaches that induce composition in diffusion planners. We find that there are two key ingredients: local receptiveness [52] during generation, and shift equivariance of states in a trajectory [36]. Through experiments in navigation environments, we show that although most current diffusion planning models do not possess this compositional capacity natively, it can be induced through architecture changes, inference choices, data augmentation strategies, or data scaling without needing dynamic programming or TD-learning. We further show that the diversity of trajectories this compositionality enables can be selectively sampled from using inpainting, allowing the use of diffusion planners only trained on imitation as goal-conditioned planners, similar to offline goal-conditioned RL.
Our primary contributions are as follows:
@@ 1 Introduction
Showing that local attention and shift equivariance, previously shown to mechanistically explain creativity and generalization in image diffusion models [36], also function as ingredients for composition in diffusion planners. However, unlike the image diffusion case, they are not the only factors: the inference method used to sample trajectories from diffusion planners also plays a significant role.
@@
Demonstrating that when induced through architecture changes, the stitching capabilities of whole-sequence diffusion can be effectively guided through simple inpainting to achieve strong goal-conditioned performance on unseen start-goal pairs.
@@
Our work provides more clarity behind what components are necessary to guide diffusion planners to compose trajectories at test-time successfully. We hope this will inspire future algorithmic and architectural developments in compositional diffusion planners.
2 Related Work
@@
Diffusion Models: Diffusion models [61, 30] offer an efficient way train models that can sample from complex distributions. Diffusion models train from data points drawn from some target distribution by taking a sample and iteratively adding Gaussian noise, such that , called the forward diffusion process. The diffusion model learns the reverse diffusion process, where some model (usually a neural network) parameterized by is trained to learn by minimizing the MSE loss between the model’s output and the reverse diffusion step: . Sampling from the new distribution is performed by first sampling from the Gaussian and using the model to reverse-sample the diffusion process: [50, 7].
@@
Diffusion Planning: Introduced by [34], diffusion planners model trajectories with diffusion. A diffusion model predicts an upcoming state or state-action trajectory horizon time-steps long, conditioned on a starting state or state memory of length , in an environment with max steps. From the generated plan, actions are executed for a fixed horizon before generating a new trajectory. In this work, we are mostly interested in the whole sequence planning domain, where , and no replanning occurs. Diffusion planning should not be confused with the diffusion policy [13] formulation where the diffusion model generates a short sequence of actions with high reactivity from frequent re-generation of actions. We elaborate on this distinction further in Appendix A.6.3.
Stitching in Imitation Learning: While stitching is often viewed as a strength of TD learning [38], other work analyzes stitching in imitation learning [8, 28]. We were inspired by [26], which analyzes stitching through a compositional generalization perspective. They revealed that applying supervised-learning as a form of offline RL [21] fails to stitch, requiring training data augmentation which manually stitches together trajectories. However, their result does not directly translate to ours because the properties we analyze (local receptiveness and shift equivariance) both enforce a form of regularization that creates composition, even when such behaviour is not explicitly built into the data.
@@
Compositionality in Diffusion Planners: Many approaches exist to create compositionality in diffusion planners. One common thread is performing hierarchical planning, including using diffusion to stitch existing dataset trajectories [41], generating subsequences jointly conditioned on each-other [48], or using a hierarchical cascade of models to separately create high-level plans and low-level trajectories [10, 42]. Our work differs from these by mainly targeting whole-sequence compositionality, where composition does not occur from explicit stitching but from implicit model behaviour, similar to image diffusion. 111See Appendix A.6.2 for more discussion.
Several other works have observed that Diffusion models often produce stitched trajectories when frequently replanning during inference [9] and limiting the history the next generation is conditioned on [62]. Our work attempts to formalize these qualitative observations with more robust experiments and explain it under a single conceptual framework (locality for short history and shift equivariance for the replanning process) that also extends to architectural or training choices.
Our focus on shift equivariance and locality was inspired by prior works which observed that these properties explain much of the compositional behaviour of image diffusion models [36, 52]. We extend their work in a few significant ways. One, we connect their image-diffusion observations to sequential decision-making by connecting these properties that enable composition to other methods commonly used to achieve stitching in IL. Two, we show that common diffusion architectures fail to exhibit these properties, preventing stitching from occurring, especially at small data sizes. Three, we analyze the effect each independently has on compositionality. Finally, we show that common guidance strategies enable strong base-model compositionality to be directed towards certain goals, which is particularly important when using generative models for offline goal-conditioned RL.
3 Methodology
3.1 Local Receptiveness and shift equivariance in Score-Based Diffusion Models
To unify prior explanations for compositionality in diffusion planners, we rely on two properties previously described for diffusion models in vision [36, 52].
The first is locality, which means the size of the region around an individual state in a trajectory that influences the next denoising step for that state is constrained to a small region. A local model only attends to a small patch around part of a sample when denoising, but through multiple denoising steps, information can propagate across a generated sample.
Definition 1: Locally Receptive Field. Assume we are interested in generating trjectories , and have a generative trajectory model whose output estimates the denoising gradient . We say that has a field locally receptive to a state neighbourhood if for all trajectories and state positions within those trajectories , .
The second is positional equivariance, which requires that the model not use the absolute position of a state in the sequence as a condition during the denoising process. Convolutional neural networks (CNNs) are typically assumed to have this property, as convolutions by design shift a shared weight kernel and apply it to local patches in a sequence. However, downsampling and pooling operators common in CNNs break this property [70, 6].
Definition 2: Shift equivariance. A trajectory score model is shift equivariant if for all shifts 222By shift equivariant we mean to discrete shifts and not continuous translations. See [49]. on trajectories which uniformly change the position of states within the trajectory, . 333In theory, circular padding is required for a CNN to have full shift equivariance[36], but in practice this is uncommon.
3.2 Composition in Diffusion Planners
Our main insight is that these notions of locality and shift equivariance have analogies to other strategies used to increase composition in diffusion planners.
Increasing Compositionality by Replanning: The first strategy we address is replanning, specifically replanning with a limited history. A trajectory planning model replans when it generates a trajectory of length , executes some action horizon , and then generates a new trajectory starting at the current state . Because a diffusion planner can be queried to generate plans with a variable length, is often set to some number large enough to ensure temporal consistency far enough into the future to do effective planning. Additionally, a history of past states to condition on with length must be decided.
Formally, given a generative trajectory model that generates trajectories of length conditioned on a history of length , each state in the trajectory will only be directly attentive to a local field of size , i.e, .
It is important to note that locality through replanning and locality in the reverse-diffusion process are not identical, as the reverse diffusion process being several steps long means influence can be spread along a long trajectory through many local denoising steps. We show this visually in Figure 3. However, the similar mechanisms highlight how increasing the influence of nearby states in a given trajectory is important for composition.
Replanning also ensures a form of shift equivariance when replanning with a fixed amount of the past trajectory kept as conditioning for the future trajectory. Diffusion planners can condition their plans on past state observations in two ways: either by including it as an explicit condition to the diffusion network, usually combined with classifier-free guidance (CFG) [31], or by inpainting states in the new generated trajectory. Full memory with inpainting does not guarantee shift equivariance, as each state in the trajectory can end up only being drawn to the same position in the trajectory that they were seen in during training. However, replanning with a smaller history shifts the viewed positions of a particular segment of the trajectory over by steps. Paradoxically, this means models with strong replanning capability must already have some form of shift equivariance.
Increasing Compositionality by Training Augmentation: The next strategy to increase compositionality we address is training augmentation, where the data is modified during training to give the model approximate learned equivariance to some transformation [3, 59]. The most common approach to encourage stitching is manually stitching together trajectories in the training set randomly by finding similar states in different trajectories[26, 8, 28, 39]. This augmentation attempts to breaks the corelation between states far away in a trajectory that prevents stitching. However, this strategy requires tuning the distance metric that determines when segments are close, and is potentially infeasible if the combinatorial set of all legal stitches is large enough to make augmenting every possible trajectory expensive.
We instead focus on augmenting the data to provide shift equivariance. Specifically, we examine positional augmentation, where the position of specific states within a trajectory are randomized during training such that the model cannot overfit to a state’s position. We implement this by sampling a random state from a training trajectory, removing the states before that state, and then padding the end of the trajectory with the final state until it is the same length as the original trajectory, such that the trajectory is randomly "shifted" by a number of states.444This assumes that the environment can be kept at the same state by the agent, which is true of our environments. Alternatives include masking out the remainder of the trajectory or allowing for mixed-length trajectories during training.
Formally, given a training trajectory with states , we sample a position from the uniform distribution . To create our new augmented trajectory , we shift the states in by positions to the left, such that for state indexes , . For the remainder of the states , we set pad the states with the last state in the original trajectory: .This modified trajectory is then used when training our planner.
Increasing Compositionality via Architecture Design Choices: The next strategy we address are architecture modifications. While equivariance can be learned through data augmentation, another popular strategy is to simply build equivariance to certain transformations through architecture changes[14, 15], especially in limited data settings.
Surprisingly, we found that the common 1D U-Net proposed in [34] has a globally receptive field and is not equivariant to shifts, which we show in Appendix A.4.1. To easily analyze the effect that locality and shift equivariance have on composition, we replace the U-Net with a network which we call Eq-Net. Eq-Net simply consists of stacked 1D convolutions without pooling or down-sampling layers, which guarantees equivariance to shifts. We also decrease the size of the kernels in Eq-Net to guarantee locality. We defer further discussion and explanation to Appendix A.4.2.
To test the effects of having a network with local receptiveness but not shift equivariance, we add positional encoding to Eq-Net with sinusoidal embeddings [66], adding each time positions’ embedding channel-wise after the first convolutional layer. See Appendix A.4.5 for details.
Increasing Compositionality via Scaling Training Data: The last strategy we address is scaling training data, which has been frequently observed to increase diffusion model generalization [68] and compositionality [53]. Importantly, this increase in compositionality does not come from memorizing more of the legal space of generations, but from learning the underlying rules governing composition in a particular area. Specifically, [23] finds that as data scales, progressively higher-level grammar rules are learned, starting with low level (local) consistency rules.
Our work is specific to sequential decision making, where data is often sparse. For instance, the toy vision dataset CIFAR-10 [37] has samples, and the more modern ImageNet [17] has . By contrast, most of the environments in the modern offline RL OGBench dataset [54] have between and episodes per environment. Similarly, most of the tasks in [13] use or less trajectories. Given that most sequential decision making datasets have orders of magnitude fewer samples, considering other approaches to generalization becomes more important than relying on data scaling.
Explaining Composition in Other Diffusion Planners: Our goal in examining the factors that lead to composition is to help explain why prior works compose (or fail to) so that future diffusion planning practitioners know which choices are important if their domain requires compositionality. Due to space constraints we include a thorough discussion of other works in Appendix A.6.3. Our main finding is that the majority of existing works achieve locality and shift equivariance either through frequent replanning with short-history conditioning [62, 9, 13] or some form of hierarchical decomposition where trajectories are generated explicitly as sub-trajectories and then composed manually [10, 48, 42]. Surprisingly, achieving compositionality through architecture choices alone is rare, with our work to our knowledge being the first to take this approach in the planning domain.
4 Experiments
Our hypothesis is that local receptiveness and positional equivariance are key ingredients in explaining compositionality in diffusion planning, and that different ways of incorporating each (training augmentations, replanning, architecture choices, etc.) should increase compositionality. However, it is not known which of the two properties is more important, whether both are needed, or which way of inducting the two properties increases compositionality the most effectively.
To test our hypothesis that local receptiveness and shift equivariance are sufficient for composition in diffusion planners and answer our open questions, we train a variety of models on toy environments designed to test generated sequence diversity and multi-goal completion.555Our code can be found here: https://github.com/rvl-lab-utoronto/diffusion-stitching
Environment Description: Our main environment is a navigation task where an agent must navigate from one point in a grid to another while avoiding obstacles. Figure 1 shows examples of our grid environment, including points navigating to and from different locations in the grid. Additionally, to test training data scaling, we use a separate navigation environment where an agent moves between ten points in euclidean space in different sequences of points. More details are deferred to Appendix A.3.
We chose toy navigation environments over complex benchmarks for a few of reasons. First, popular datasets which claim to test stitching either have testing start-goal pairs contained in the training set [24]666Shown by [26]. or a very limited number of start and goal states [54]. Additionally, existing datasets do not report what subsequences are present in the training data, making it difficult to predict stitching behaviour, whereas we can control this in our environment. Finally, the grid layout of our environment makes quantifying diversity easy, as each stitched trajectory will be topologically distinct [57], allowing us to perform careful analysis of a given model’s compositional ability.
Data Collection for Diffusion Model Training: We look at the diversity of trajectories generated by diffusion planners in two separate planning settings, and collect two seperate datasets in our grid environment.
The first is the fixed-endpoint setting, where the goal of the task is to reach a fixed end point from a fixed start point. Every training trajectory is optimal for the task, but the demonstration data includes multi-modal behaviour, as individual trajectories reach the end point via a different route. All trajectories have the same starting and ending state but take different paths through the grid, and the diffusion planner is tasked with generating new paths from the start to the goal composed of sub-trajectories seen in training in new configurations.
The second is the end-point conditional setting, where a wider range of trajectories provide sub-trajectories that need to be composed to meet novel start-goal pairs. In our environment, the goal is defined as reaching a goal state from a start state, where each intersection point along the edge of the grid is a legitimate start or goal state. To be precise, we attempt to learn a model that can sample from the distribution . We collect data that includes diverse grid navigation to give the model ample sub-trajectories to compose, while being careful to keep some start-goal pairs unseen. We defer details to Appendix A.3.
Algorithm Setup: For our experiments we use a 1D U-Net Diffusion Planner as the base model. We use CleanDiffuser’s [18] implementation and hyperparameters. One change we make is only predicting future states and using an inverse dynamics model to recover actions [1]. 777See Table 2 for details on hyperparameters and inverse dynamics..
5 Results
Here we report our findings on the unconditional and conditional environments, with the environments, data, and algorithms described in Section 4.
In our results, we count diversity by examining topologically distinct trajectories. Practically, each trajectory can split the obstacles into two groups: those on its left side, and on its right; two trajectories with different obstacle splits are topologically unique.
5.1 Unconditional Environment Results
To quantify each model’s ability to compose in the unconditional environment, we took our final trained models and sampled them at a range of sample counts. From these samples, we counted what percentage of the possible (but unseen in training) trajectory space each model’s generations covered. By giving us a sense of how a model’s coverage scales with the number of generations, this gives a sense of how capable each model’s composition abilities are without solely relying on a model’s behaviour at very high sampling counts.888In our unconditional results, we filter to only analyze successful trajectories (ones that successfully reached the end state without hitting an obstacle) but all methods achieved very high success rates. Full completion rates are reported in Appendix A.2.1.
Figure 4 shows our results for the unconditional navigation environment. Figure 4(a) shows a comparison of the different whole-sequence composition strategies. The base U-Net fails to ever compose on our small dataset, only reproducing examples seen in training. Our model with approximate positional equivariance (granted through training augmentation) performs better, but fails to extrapolate, not covering a significant proportion of the unseen trajectory space even at very high sampling rates. Our equivariant network with added positional encoding (making the model local, but not positionally equivariant) performs much better, indicating locality is a more important factor in encouraging composition than positional equivariance. Lastly, our model that is both locally receptive and shift equivariant (Eq-Net) performs the best, exhibiting very strong whole-sequence composition from architecture choices alone, without utilizing data scaling, augmentation, or replanning.
Figure 4(b) analyzes replanning, where we progressively increase the number of times the model replans during a trajectory, and analyze the resulting diversity of generations over independent generations. For our replanning model we used our U-Net trained with positional data augmentation, as we found it necessary to enable the model to generate trajectories from diverse start positions. Assume we wish to replan times in a task long. When replanning, we keep only the current state ( using our prior notation), generate a new trajectory starting at that point, execute steps, and then repeat until the goal is reached. Figure 4(b) shows that as expected, as , the diversity of generations increases. To compare this to our architectural strategy, we compare it against the best model from Figure 4. To directly compare the computational cost associated with replanning, we compare each agent that replans times against Eq-Net that gets times as many parallel samples to generate from, such that each point queries the diffusion model an equal amount of times. For instance, at , the replanning model over independent samples is queried times, as each trajectory requires queries, so to have a fair comparison we query Eq-Net for independent samples.
Predictably, our results show that more frequent replanning indeed leads to more trajectory diversity. Interestingly, they also show that the extra budget of diffusion model queries required during replanning might be better spent generating more samples from a high-diversity whole-trajectory model, as Eq-Net performs similarly or slightly better on all samples.
Lastly, to validate that our results are not influenced by regularization induced by under-training of our diffusion planner, we analyze how each model’s compositionality evolves throughout training. We find that while early stopping helps compositionality (perhaps acting as a form of regularization that prevents overfitting to the training samples as has been observed in other work [5]), models with stronger inductive biases preserve their compositional abilities as they trained for longer. See Appendix A.5.1 for details.
5.2 Conditional Environment Results
In the conditional environment described in Section 4, the goal of the diffusion planner is to take unstructured trajectories in our environment and be able to consistently produce trajectories connecting an arbitrary start and goal location. To ensure that stitching is required at inference time, we examine every pair of edge intersections and filter out any pair where the starting pair and goal pair appear together sequentially, resulting in test-time pairs. We run our whole-sequence planners on each pair times, calculate an average completion rate and number of unique successful plans (where uniqueness is defined as topological distinctness), and take the average over sample runs.
To guide the model to a specific goal, we use inpainting like in prior diffusion works [47, 34, 61]. We found that repeating the inpainted goal state significantly improved goal-reaching performance, so we repeat the goal state times at the end of the trajectory before diffusing the intermediate states.
Our results are shown in Figure 5. The model that is both local and positionally equivariant far outperforms all other models, achieving an average completion rate, roughly higher than the model with only locality but not positional equivariance. Only having positional equivariance did not substantially help the model in this scenario, with the U-Net with data augmentation only achieving a non-zero result on one of the testing pairs and only a small amount of the time. Lastly, like in the unconditional case, the default U-Net fails to perform stitching entirely if it is not locally receptive or shift equivariant.
To visualize the difference in composition for each model, we plotted the unique successful runs from random generations from each algorithm on one of the hold-out start-goal pairs. Figure 6 shows the relative diversity of each model: the base U-Net fails to ever generate a successful trajectory, the U-Net made approximately positionally equivariant with data augmentation can produce a single candidate successful trajectory, the Eq-Net with positional equivariance removed produces a handful of plans, and the full Eq-Net produces a large diversity of successful trajectories.
5.3 Data Scaling Results
Lastly, we investigate the role of data in inducing composition. We use a different environment in these experiments, consisting of a navigation task where a point must sequentially navigate between five different points of a decagon. A compositional planner should be able to stitch sub-trajectories from individual points to other points. We train models with data ranging from from to data points, where points is close to the number of possible combinations.
Figure 7 shows that even models without strong inductive biases can learn to compose at large enough data sizes. Past a certain data quantity (roughly samples in our environment), the U-Net learns to compose, producing new stitched trajectories at test time. We provide evidence that this data scaling induces a smaller receptive field of the diffusion model in Appendix A.5.2. This confirms that the ability to generalize composition of sub-sequences is another property that diffusion models learn from data at scale, adding onto concept composition [53], consistency [68] and grammar [23].
6 Discussion
6.1 What Are the Keys to Composition?
While our work attempts explain why existing works in diffusion planning achieve composition, a reasonable question to ask is what engineering insights our findings have. While our results indicate that achieving whole-sequence compositionality through architecture choices may be under-explored, we emphasize that different choices are valid in different contexts. For instance, in environments where reactive control is very effective, generating long-horizon sequences may not be important, so frequent replanning with a short horizon may work better. In environments where data is abundant, using a higher-capacity model might be appropriate. We provide additional discussion of our findings in Appendix A.6.
6.2 Limitations and Future Work
Our work has several limitations. First, we only consider navigation environments with little complexity in the dynamics or observation space. While we motivate this in Section 5, evaluating our findings in a more realistic robotics environment would be useful future work. Second, our guidance strategy was limited to inpainting, which is effective in environments with clear endpoint goals, but less useful when goals are abstract (such as with language-conditioned agents) or multi-modal. Incorporating a way to leverage other guidance strategies like classifier-based [30] or free [31] guidance into our proposed high-compositionality diffusion planners is potentially fruitful future work. This would also let us fit our composition framework into the diffusion policy [13] setting, where composition of atomic skills may still be useful, but where a goal state cannot be inpainted because the diffusion model generates future actions and not future states.
Appendix A Appendix
Contents
A.1 More Related Work
Prior to diffusion planners, other generative models were common in imitation learning [29]. Most modern approaches are either based around autoregressive decoder-only transformers [58], such as the Decision Transformer [12], ACT [25], or diffusion models [29]. A more comprehensive survey on Diffusion Planners was prepared by [63].

### v3

# What Do You Need for Compositional Generalization in Diffusion Planning?

@@ Abstract
In policy learning, stitching and compositional generalization refer to the extent to which the policy is able to piece together sub-trajectories of data it is trained on to generate new and diverse behaviours. While stitching has been identified as a significant strength of offline reinforcement learning, recent generative behavioural cloning (BC) methods have also shown proficiency at stitching. However, the main factors behind this are poorly understood, hindering the development of new algorithms that can reliably stitch by design. Focusing on diffusion planners trained via generative behavioural cloning, and without resorting to dynamic programming or TD-learning, we find three properties are key enablers for composition: shift equivariance, local receptive fields, and inference choices. We use these properties to explain architecture, data, and inference choices in existing generative BC methods based on diffusion planning including replanning frequency, data augmentation, and data scaling. Our experiments show that while local receptive fields are more important than shift equivariance in creating a diffusion planner capable of composition, both are crucial. Using findings from our experiments, we develop a new architecture for diffusion planners called Eq-Net, that is simple, produces diverse trajectories competitive with more computationally expensive methods such as replanning or scaling data, and can be guided to enable generalization in goal-conditioned settings. We show that Eq-Net exhibits significant compositional generalization in a variety of navigation and manipulation tasks designed to test planning diversity.
I Introduction
The advent of diffusion models [36] has inspired a wave of work using these models as planners and policies in robotic systems [40, 13]. Image and video diffusion models exhibit strong compositionality: the ability to take component elements seen in samples of their dataset and compose them in new, feasible ways [37, 21, 52]. When and why diffusion models for planning and control compose, however, is not well understood. Diffusion Policies [13], which generate short action chunks, naturally compose behaviours through frequent replanning. However, Diffusion Planners [40], which generate long-horizon plans, often struggle with composition [28, 9], instead reproducing exact training trajectories.
Composition in diffusion planners is crucial to address, as the capability to compose behaviours into new trajectories at test time is important property for planners broadly. Reasons include allowing for better candidate set selection [79], encouraging multi-modal behaviour [47], and generating better data for offline reinforcement learning [45].
To address this gap, we analyze a variety of approaches that induce composition in diffusion planners. We find that there are two key ingredients: the ability to use sub-skills at any point in a sequence (shift equivariance [42]), and a locality bias towards attending to nearby states when generating a state in a trajectory (local receptive fields [62]). We show that previous methods achieve composition through these two ingredients, identifying composition can be ensured through careful inference choices, data augmentation strategies, or data scaling.
Using these two ingredients, we develop a neural network architecture for the diffusion planning denoising network called Eq-Net, which is simple, produces diverse trajectories, and can be guided to enable strong goal-conditioning. We show Eq-Net produces more diverse trajectories than other architectures in a variety of navigation and robotic manipulation settings.
Our contributions are as follows:
@@ I Introduction
We show that local receptive fields and shift equivariance are key enablers for composition in diffusion planners, as has been shown in the image generation domain [62, 42]. However, unlike the image generation case, they are not the only important factors: the inference method used to sample trajectories from diffusion planners also plays a crucial role in achieving compositional generalization.
We show that stitching achieved through data scaling, as opposed to architectural choices, still induces local receptive fields in models.
@@
We demonstrate that when induced through architecture changes, the stitching capabilities of whole-sequence diffusion can be effectively guided through inpainting to achieve strong goal-conditioned performance on unseen tasks.
@@
Our work aims to provide clarity behind the key factors that guide diffusion planners to compose trajectories at test-time, informing future algorithmic and architectural developments in compositional diffusion planners.
II Preliminaries and Related Work
@@
Diffusion Models: Diffusion models [71, 36] are generative models trained on data points drawn from some target distribution by taking a sample and iteratively adding Gaussian noise, such that , called the forward diffusion process. The diffusion model learns the reverse diffusion process, where a neural network is trained to learn by minimizing the MSE loss between the model’s output and the reverse diffusion step: . Sampling is performed by sampling from the Gaussian and using the model to reverse-sample the diffusion process by taking repeated steps of size from : [60, 7]. Many different formulations to diffusion model prediction exist, but most works either predict a clean sample from noise [49] or the residual noise to remove [36].
@@
Diffusion Planning: Introduced by Janner et al. [40], diffusion planners model entire trajectories by sampling from diffusion models. A diffusion planner generates an upcoming state trajectory consisting of a sequence of states horizon time-steps long, sometimes conditioned on a state memory of length in an environment with max steps. Actions from the plan are executed for a horizon before generating a new trajectory. Here we mostly investigate the whole-sequence planning domain, where , and no replanning occurs. The more common Diffusion Policy [13] formulation, where the diffusion model generates a short sequence of actions with high reactivity from frequent re-generation of actions, can be thought of as a special form of diffusion planning where . We elaborate on this distinction further in Appendix VI-H3.
@@
Compositionality in Diffusion Planners: Many approaches exist to explicitly induce compositionality in diffusion planners. Hierarchical approaches use diffusion to stitch existing dataset trajectories [48], generate subsequences jointly conditioned on each-other [57], or separately generate high and low level plans[10, 50]. By contrast, we target whole-sequence compositionality, where composition does not occur from explicit stitching but from implicit model behaviour, similar to image diffusion. Benefits of a monolithic planning approach includes reducing algorithmic complexity, sidestepping potentially harmful planning hierarchy [2], and enhanced performance in long-horizon tasks [9, 54].